If a random variable u has t -distribution with n degree of freedom, find the 

distribution of u2.


1
Expert's answer
2021-06-07T19:03:42-0400

Solution:

Given, "u" has t-distribution.

So, "u\sim \dfrac{Z}{\sqrt{X/n}}=\dfrac{N(0,1)}{\chi^2_n/n}"

Then, "u^2\sim \dfrac{Z^2}{{X/n}}=\dfrac{\chi^2_1/1}{\chi^2_n/n}=\dfrac{n\chi^2_1}{\chi^2_n}"

Thus, distribution of "u^2" is "\dfrac{n\chi^2_1}{\chi^2_n}"